# Dominated Actions in Imperfect-Information Games

### v6

@@ 4 Algorithm for identifying dominated actions in two-player games
Suppose we want to determine whether an action $c$ for player 1 is dominated at information set $I$ in a two-player perfect-recall extensive-form game $G$. Using the sequence form representation, suppose that action $c$ is the final action in the sequence with index $i$ for player 1. Let $S_{I}$ denote the set of indices corresponding to action sequences leading to $I$. We would like to solve the following problem.
@@
Let $f_{1}$ be the optimal objective value in Problem 1 and $f_{2}$ be the optimal objective value in Problem 2. Suppose that the optimal variables in Problem 1 are $\mathbf{x}^{1}_{1},\mathbf{x}^{1}_{2},\mathbf{y}^{1}.$ Now set $\mathbf{x}^{2}_{1}=\mathbf{x}^{1}_{1},$ $\mathbf{x}^{2}_{2}=\mathbf{x}^{1}_{2},$ $\mathbf{y}^{2}_{1}=\mathbf{y}^{1},$ $\mathbf{y}^{2}_{2}=\mathbf{y}^{1}.$ Then $(\mathbf{x}^{2}_{1},\mathbf{x}^{2}_{2},\mathbf{y}^{2}_{1},\mathbf{y}^{2}_{2})$ gives a feasible solution to Problem 2 with objective value equal to $f_{1}.$ So $f_{2}\geq f_{1}$. Now suppose that the optimal variables in Problem 2 are $\mathbf{x}^{2}_{1},\mathbf{x}^{2}_{2},\mathbf{y}^{2}_{1},\mathbf{y}^{2}_{2}.$ Now set $\mathbf{x}^{1}_{1}=\mathbf{x}^{2}_{1},$ $\mathbf{x}^{1}_{2}=\mathbf{x}^{2}_{2},$ and set $\mathbf{y}^{1}$ equal to the strategy that follows $\mathbf{y}^{2}_{1}$ at states following action $c$ for player 1, and follows $\mathbf{y}^{2}_{2}$ otherwise. Then $(\mathbf{x}^{1}_{1})^{T}\mathbf{A}\mathbf{y}^{1}=(\mathbf{x}^{1}_{1})^{T}\mathbf{A}\mathbf{y}^{2}_{1},$ since both players only take actions to get to information set $I$, $\mathbf{x}^{1}_{1}$ takes action $c$ at $I$, and the strategies $\mathbf{y}^{1}$ and $\mathbf{y}^{2}_{1}$ are identical after player 1 takes action $c$ at $I.$ Similarly, $(\mathbf{x}^{1}_{2})^{T}\mathbf{A}\mathbf{y}^{1}=(\mathbf{x}^{1}_{2})^{T}\mathbf{A}\mathbf{y}^{2}_{2},$ since both players only take actions to get to $I$, $\mathbf{x}^{1}_{1}$ does not take action $c$ at $I$, and the strategies $\mathbf{y}^{1}$ and $\mathbf{y}^{2}_{2}$ are identical after player 1 does not take action $c$ at $I.$ So $f_{1}\geq f_{2}$. So we conclude that $f_{1}=f_{2}.$
Proposition 1 allows us to divide Problem 2 into the following two subproblems. If $f$ is the optimal objective value of Problem 2, $f_{1}$ is the optimal objective of Problem 3, and $f_{2}$ is the optimal objective of Problem 4, then we have $f=f_{1}-f_{2}.$
@@
$L(\mathbf{y},\boldsymbol{\lambda},\boldsymbol{\gamma},\mathbf{r})=\mathbf{x}^{T}_{2}\mathbf{A}\mathbf{y}-(\mathbf{f}^{T}-\mathbf{y}^{T}\mathbf{F}^{T})\boldsymbol{\lambda}-\sum_{k\in S_{I}}\boldsymbol{\gamma}_{k}(\mathbf{y}_{k}-1)-\mathbf{r}^{T}\mathbf{y}$
$\frac{\partial L}{\partial\mathbf{y}}=\mathbf{x}^{T}_{2}\mathbf{A}+\boldsymbol{\lambda}^{T}\mathbf{F}-\sum_{k\in S_{I}}\boldsymbol{\gamma}_{k}\mathbf{e_{k}}-\mathbf{r}^{T}$
@@
Next consider Problem 4. Note that both players are aligned in trying to maximize the objective. Let us define a new problem $\overline{G}$ where player 1 selects the actions for both players. We denote player 1’s strategy in this modified game by $\overline{\mathbf{x}}_{1}.$ We modify $\mathbf{E}$ and resize $\mathbf{e}$ accordingly and denote them by $\mathbf{E^{\prime}}$ and $\mathbf{e^{\prime}}.$ The payoffs can now be represented as a vector $\mathbf{a}$. In the new representation let $i^{\prime}$ denote the index of the sequence with concluding action $c$, and let $S^{\prime}_{I}$ denote the set of indices corresponding to action sequences leading to $I$.
@@
Let $u_{2}$ denote the optimal objective value for optimization problem 5, and $u_{1}$ denote the optimal objective value for problem 6. If $u_{2}>u_{1}$ then we conclude that action $c$ is strictly dominated. If $u_{2}<u_{1}$ then we conclude that action $c$ is not strictly or weakly dominated. If $u_{2}=u_{1}$, let $u_{3}$ denote the optimal objective value for problem 7, and let $u_{4}$ denote the optimal objective value for problem 8. If $u_{3}>u_{4}$ then we conclude that action $c$ is weakly dominated, and if $u_{3}=u_{4}$ we conclude that action $c$ is not strictly or weakly dominated (note that we cannot have $u_{3}<u_{4}$).
@@
We can perform this procedure for every action at each information set of player 1, and analogously for player 2. Since the number of actions is linear in the size of the game tree, the overall procedure involves solving a linear number of linear programs and therefore runs in polynomial time. We can repeat the procedure iteratively until no more actions are removed for any player. Thus, iterative removal of dominated actions can be performed in polynomial time. Note that the procedure applies to all two-player games and does not assume that they are zero sum. The procedure also removes actions that are dominated by any mixed strategy (which may play a probability distribution over actions at the same information set $I$), not just actions that are dominated by a pure action.
@@
There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in a two-player perfect-recall extensive-form game of imperfect information.
@@
There exists a polynomial-time algorithm that determines whether a given action is weakly dominated in a two-player perfect-recall extensive-form game of imperfect information.
@@
Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in a two-player perfect-recall extensive-form game of imperfect information.
5 Extension to $n$ players
Suppose we wish to determine whether an action of player 1 is strictly or weakly dominated in an $n$-player perfect-recall game $\Gamma$ with $n>2$. Although the linear-programming characterization developed above applies only to two-player games with perfect recall, we can reduce the general case to the two-player setting by merging all players other than player 1 into a single aggregated opponent. This merged player controls every decision node that originally belonged to players $2,\dots,n$, and its information sets are refined as needed to preserve perfect recall. The resulting game $\Gamma^{\prime}$ has exactly two players (player 1 and the merged opponent), while preserving the game tree, player 1’s information partitions, and player 1’s payoffs.
A key observation is that strict (or weak) dominance for player 1 is preserved under this transformation: an action is dominated in the original $n$-player game if and only if it is dominated in the merged two-player game. Therefore, once the game is transformed into $\Gamma^{\prime}$, we may directly apply the polynomial-time dominance tests developed above for the two-player case. In this way, both single-step dominance testing and iterated removal of dominated actions can be performed in polynomial time even in general $n$-player perfect-recall extensive-form games.
Let $\Gamma$ be a finite $n$-player extensive-form game of imperfect information with perfect recall, with players $N=\{1,2,\dots,n\}$. Fix player 1 and an information set $I_{1}$ of player 1. We construct a two-player game $\Gamma^{\prime}$ with players $1$ and $P$ (the “merged” opponent) as follows.
  • The underlying game tree (histories, chance nodes, terminal nodes, and payoffs) of $\Gamma^{\prime}$ is identical to that of $\Gamma$. In particular, the payoff function $u_{1}$ for player 1 is unchanged.
  • Every decision node controlled by player 1 in $\Gamma$ is still controlled by player 1 in $\Gamma^{\prime}$. Every decision node controlled by any player $j\in\{2,\dots,n\}$ in $\Gamma$ is controlled by player $P$ in $\Gamma^{\prime}$.
  • The information sets of player 1 in $\Gamma^{\prime}$ are exactly the same as in $\Gamma$.
  • The information sets of player $P$ are obtained from the information sets of players $2,\dots,n$ in $\Gamma$ by refinement to ensure perfect recall for $P$: for each original information set $J$ of some player $j\in\{2,\dots,n\}$, we partition $J$ into the equivalence classes induced by the relation $h\sim h^{\prime}$ if and only if the sequence of information sets and actions at which any of the original opponents $\{2,\dots,n\}$ moved along the path to $h$ is identical to that along the path to $h^{\prime}$.
By construction, $\Gamma^{\prime}$ is a two-player extensive-form game with perfect recall (for players $1$ and $P$), and its size is linear in the size of $\Gamma$.
Lemma 1.
Let $\Gamma$ be an $n$-player extensive-form game of imperfect information with perfect recall, and let $\Gamma^{\prime}$ be the two-player game obtained by merging players $2,\dots,n$ into a single opponent $P$ as above. Fix an information set $I_{1}$ of player 1 and an action $a\in A(I_{1})$. Then $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ (according to Definition 1) if and only if it is strictly dominated for player 1 at $I_{1}$ in $\Gamma^{\prime}$.
Proof.
Let $\Sigma_{j}$ denote the set of behavior strategies of player $j$ in $\Gamma$, and let
$\Sigma_{-1}\;=\;\Sigma_{2}\times\cdots\times\Sigma_{n}$
denote the set of behavior-strategy profiles of all players other than 1. For an information set $I_{1}$ of player 1, let $\Sigma_{-1}^{I_{1}}\subseteq\Sigma_{-1}$ denote the subset of profiles that “always play to reach $I_{1}$ when possible,” as in Definition 1. Similarly, let $\Sigma_{P}$ denote the set of behavior strategies of player $P$ in $\Gamma^{\prime}$, and let $\Sigma_{P}^{I_{1}}\subseteq\Sigma_{P}$ be the subset that always plays to reach $I_{1}$ when possible in $\Gamma^{\prime}$.
Because behavior strategies assign an independent probability distribution to each information set, and because the information sets of $P$ in $\Gamma^{\prime}$ are obtained by refining and relabeling the information sets of players $2,\dots,n$ in $\Gamma$, there is a natural bijection
$\Phi:\Sigma_{-1}\;\longrightarrow\;\Sigma_{P}$
that maps a profile $(\sigma_{2},\dots,\sigma_{n})$ to a single behavior strategy $\sigma_{P}$ which, at each information set of some original player $j\geq 2$, plays according to $\sigma_{j}$. The local “always reach $I_{1}$” constraints are preserved under this mapping, so
$\Phi\big(\Sigma_{-1}^{I_{1}}\big)=\Sigma_{P}^{I_{1}}.$
For any profile $(\sigma_{1},\sigma_{-1})$ in $\Gamma$, and the corresponding $(\sigma_{1},\sigma_{P})$ in $\Gamma^{\prime}$ with $\sigma_{P}=\Phi(\sigma_{-1})$, the induced distribution over terminal histories is the same in $\Gamma$ and $\Gamma^{\prime}$, since the game tree and player 1’s actions are unchanged and the opponents’ local behavior at each information set is identical. Therefore, for all such profiles we have
$u_{1}(\sigma_{1},\sigma_{-1})\;=\;u_{1}(\sigma_{1},\sigma_{P}),$
where $u_{1}$ is player 1’s payoff.
Now fix an action $a\in A(I_{1})$. By Definition 1, $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ if and only if there exists a behavior strategy $\sigma_{1}^{-a}$ for player 1 that always reaches $I_{1}$ and never plays $a$ at $I_{1}$ such that, for every behavior strategy $\sigma_{1}^{a}$ that always reaches $I_{1}$ and plays $a$ with probability 1 at $I_{1}$, and for every $\sigma_{-1}\in\Sigma_{-1}^{I_{1}}$,
$u_{1}(\sigma_{1}^{-a},\sigma_{-1})\;>\;u_{1}(\sigma_{1}^{a},\sigma_{-1}).$
Using the bijection $\Phi$ and equality of payoffs, this is equivalent to: there exists $\sigma_{1}^{-a}$ such that for every $\sigma_{1}^{a}$ and for every $\sigma_{P}\in\Sigma_{P}^{I_{1}}$,
$u_{1}(\sigma_{1}^{-a},\sigma_{P})\;>\;u_{1}(\sigma_{1}^{a},\sigma_{P}),$
which is precisely the strict-dominance condition for action $a$ at $I_{1}$ in the merged two-player game $\Gamma^{\prime}$ (again using Definition 1 with the opponents collapsed into $P$).
Thus $a$ is strictly dominated in $\Gamma$ if and only if it is strictly dominated in $\Gamma^{\prime}$. ∎
Theorem 4.
There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in an $n$-player perfect-recall extensive-form game of imperfect information.
Proof.
Let $\Gamma$ be an $n$-player extensive-form game with perfect recall, let $I_{1}$ be an information set of player 1, and let $a\in A(I_{1})$ be the action to be tested.
First, construct the merged two-player game $\Gamma^{\prime}$ with players $1$ and $P$ as described above. This transformation does not duplicate any nodes of the game tree; it only relabels decision nodes and refines some information sets of the opponents. Hence the size of $\Gamma^{\prime}$ (number of nodes, information sets, and actions) is linear in the size of $\Gamma$, and the construction can be performed in polynomial time.
By Lemma 1, action $a$ is strictly dominated for player 1 at $I_{1}$ in $\Gamma$ (according to Definition 1) if and only if it is strictly dominated for player 1 at $I_{1}$ in $\Gamma^{\prime}$.
The game $\Gamma^{\prime}$ is a two-player extensive-form game with perfect recall. By the two-player result stated earlier, strict dominance of a given action in a two-player perfect-recall extensive-form game of imperfect information can be decided in time polynomial in the size of the game, using the sequence-form linear programs derived in the previous section.
Therefore, we can decide whether $a$ is strictly dominated in $\Gamma$ by:
  1. 1.
    constructing the merged game $\Gamma^{\prime}$ in polynomial time, and
  2. 2.
    running the two-player strict-dominance algorithm on $\Gamma^{\prime}$ in polynomial time.
The total running time is polynomial in the size of $\Gamma$, which proves the theorem. ∎
Theorem 5.
There exists a polynomial-time algorithm that determines whether a given action is weakly dominated for player 1 in an $n$-player perfect-recall extensive-form game of imperfect information.
Proof.
The weak-dominance test for two-player perfect-recall games is polynomial-time (Theorem 2). Weak dominance is preserved under the merging transformation exactly as in Lemma 1: the definition only modifies the payoff inequality from strict to weak, but the induced distributions and opponent strategy sets remain in bijection under the merge. Thus an action is weakly dominated in the original $n$-player game if and only if it is weakly dominated in the merged two-player game. Applying the two-player weak-dominance test to the merged game gives a polynomial-time algorithm for the $n$-player case. ∎
Theorem 6.
Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in an $n$-player perfect-recall extensive-form game of imperfect information.
Proof.
At each iteration of the iterated-removal process in the $n$-player game $\Gamma$, for every information set $I_{1}$ of player 1 and every action $a\in A(I_{1})$, we test whether $a$ is strictly or weakly dominated according to Definitions 1 and 2. To perform this test, we construct the merged two-player game $\Gamma^{\prime}$ and apply the corresponding strict or weak-dominance algorithm for two-player games.
By Theorems 4 and 5, each dominance test in $\Gamma^{\prime}$ runs in time polynomial in the size of $\Gamma$. The number of actions in $\Gamma$ is polynomial in the size of the game tree, so in each iteration we perform only polynomially many dominance tests, and each such iteration therefore runs in polynomial time. Furthermore, each action can be removed at most once, so the total number of iterations is at most the total number of actions, and hence polynomial in the size of $\Gamma$.
Finally, Lemma 1 ensures that at every iteration the dominance status of each action is identical in the original $n$-player game $\Gamma$ and in the merged two-player game $\Gamma^{\prime}$. Therefore an action is removed in $\Gamma$ at a given iteration if and only if the corresponding action is removed in $\Gamma^{\prime}$. Combining these observations, we obtain that the entire iterated-removal process runs in polynomial time and yields the same reduced game whether it is viewed in $\Gamma$ or in $\Gamma^{\prime}$. ∎

### v7

@@ 4 Algorithm for identifying dominated actions in two-player games
Suppose we want to determine whether an action $c$ for player 1 is dominated at information set $I$ in a two-player perfect-recall extensive-form game with publicly observable actions $G$. Using the sequence form representation, suppose that action $c$ is the final action in the sequence with index $i_{c}$ for player 1, and suppose that $i_{p}$ is the index of the parent sequence to sequence $i_{c}.$ Let $S_{I}$ denote the set of opponent sequences that appear on at least one path from the root to a node in information set $I.$ We restrict the opponent strategy vector so that only sequences in $S_{I}$ can receive positive probability. Each sequence corresponds to a component of the realization vector; we refer to the component index associated with sequences $s$ as its index. To consider whether action $c$ is dominated, we consider the following sequence-form optimization problem.
@@
For a fixed feasible $\mathbf{x}_{2}$, let
$\phi_{1}(\mathbf{x}_{2})=\min_{\mathbf{x}_{1},\mathbf{y}}\left(\mathbf{x}_{2}^{T}A\mathbf{y}-\mathbf{x}_{1}^{T}A\mathbf{y}\right)$
denote the inner objective value of Problem 1, and let
$\phi_{2}(\mathbf{x}_{2})=\min_{\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2}}\left(\mathbf{x}_{2}^{T}A\mathbf{y}_{2}-\mathbf{x}_{1}^{T}A\mathbf{y}_{1}\right)$
denote the inner objective value of Problem 2.
First observe that $\phi_{2}(\mathbf{x}_{2})\leq\phi_{1}(\mathbf{x}_{2})$ for every feasible $\mathbf{x}_{2}$, since any feasible pair $(\mathbf{x}_{1},\mathbf{y})$ for Problem 1 yields a feasible triple $(\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2})$ for Problem 2 by setting $\mathbf{y}_{1}=\mathbf{y}_{2}=\mathbf{y}$.
Now fix a feasible $\mathbf{x}_{2}$ and let $(\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2})$ be any feasible solution of the inner minimization in Problem 2. Define a strategy $\mathbf{y}$ that follows $\mathbf{y}_{1}$ at all information sets reachable after player 1 takes action $c$ at information set $I$, and follows $\mathbf{y}_{2}$ otherwise.
Here we use the assumption that actions are publicly observable. Because player 1’s action at $I$ is publicly observed, no information set of player 2 can contain nodes reached both after action $c$ and after a different action at $I$. Therefore, the prescriptions of $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ are never assigned inconsistently at the same information set, so this stitching defines a valid behavioral strategy $\mathbf{y}$ for Problem 1. Moreover,
$\mathbf{x}_{1}^{T}A\mathbf{y}=\mathbf{x}_{1}^{T}A\mathbf{y}_{1},\qquad\mathbf{x}_{2}^{T}A\mathbf{y}=\mathbf{x}_{2}^{T}A\mathbf{y}_{2},$
because $\mathbf{x}_{1}$ takes action $c$ at $I$ while $\mathbf{x}_{2}$ does not. Hence
$\mathbf{x}_{2}^{T}A\mathbf{y}-\mathbf{x}_{1}^{T}A\mathbf{y}=\mathbf{x}_{2}^{T}A\mathbf{y}_{2}-\mathbf{x}_{1}^{T}A\mathbf{y}_{1}.$
Since this holds for every feasible $(\mathbf{x}_{1},\mathbf{y}_{1},\mathbf{y}_{2})$, we obtain $\phi_{1}(\mathbf{x}_{2})\leq\phi_{2}(\mathbf{x}_{2})$. Thus $\phi_{1}(\mathbf{x}_{2})=\phi_{2}(\mathbf{x}_{2})$ for every feasible $\mathbf{x}_{2}$. Maximizing over $\mathbf{x}_{2}$ in the two problems yields $v_{1}=v_{2}$. ∎
Proposition 1 allows us to divide Problem 2 into the following two subproblems. If $v_{2}$ is the optimal objective value of Problem 2, $v_{3}$ is the optimal objective of Problem 3, and $v_{4}$ is the optimal objective of Problem 4, then we have $v_{2}=v_{3}-v_{4}.$
@@
$L(\mathbf{y},\boldsymbol{\lambda},\boldsymbol{\gamma},\mathbf{r})=\mathbf{x}^{T}_{2}\mathbf{A}\mathbf{y}-(\mathbf{f}^{T}-\mathbf{y}^{T}\mathbf{F}^{T})\boldsymbol{\lambda}-\sum_{k\notin S_{I}}\boldsymbol{\gamma}_{k}\mathbf{y}_{k}-\mathbf{r}^{T}\mathbf{y}$
$\frac{\partial L}{\partial\mathbf{y}}=\mathbf{A}^{T}\mathbf{x}_{2}+\mathbf{F}^{T}\boldsymbol{\lambda}-\sum_{k\notin S_{I}}\boldsymbol{\gamma}_{k}\mathbf{e_{k}}-\mathbf{r}$
@@
Next consider Problem 4. Note that both players are aligned in trying to maximize the objective. Let us define a new problem $G^{\prime}$ where player 1 selects the actions for both players. We denote player 1’s strategy in this modified game by $\mathbf{x}^{\prime}_{1}.$ We modify $\mathbf{E}$ and resize $\mathbf{e}$ accordingly and denote them by $\mathbf{E^{\prime}}$ and $\mathbf{e^{\prime}}.$ The payoffs can now be represented as a vector $\mathbf{a}^{\prime}$. In the new representation let $i^{\prime}_{c}$ denote the index of the sequence with concluding action $c$, and let $i^{\prime}_{p}$ denote the index of the parent sequence.
@@
Note that $G^{\prime}$ is not treated as a two-player game and no equilibrium computation is performed. Instead, player 1 controls all decision nodes and the problem becomes a single-agent optimization problem over the game tree. The sequence-form constraints $E^{\prime}$ simply enforce valid realization plans in the tree and do not rely on perfect recall. Therefore the formulation remains valid even though merging the players may introduce imperfect recall. The same interpretation applies to Problems 7 and 8.
Let $v_{5}$ denote the optimal objective value for optimization problem 5, and $v_{6}$ denote the optimal objective value for problem 6. If $v_{5}>v_{6}$ then we conclude that action $c$ is strictly dominated. If $v_{5}<v_{6}$ then we conclude that action $c$ is not strictly or weakly dominated. If $v_{5}=v_{6}$, let $v_{7}$ denote the optimal objective value for problem 7, and let $v_{8}$ denote the optimal objective value for problem 8. If $v_{7}>v_{8}$ then we conclude that action $c$ is weakly dominated, and if $v_{7}=v_{8}$ we conclude that action $c$ is not strictly or weakly dominated (note that we cannot have $v_{7}<v_{8}$).
@@
We can perform this procedure for every action at each information set of player 1, and analogously for player 2. The procedure is summarized in Algorithm 1. Since the number of actions is linear in the size of the game tree, the overall procedure involves solving a linear number of linear programs and therefore runs in polynomial time. We can repeat the procedure iteratively until no more actions are removed for any player. Thus, iterative removal of dominated actions can be performed in polynomial time. Note that the procedure applies to all two-player games and does not assume that they are zero sum. The procedure also removes actions that are dominated by any behavioral strategy (which may play a probability distribution over actions at the same information set $I$), not just actions that are dominated by a pure action.
Algorithm 1 Dominance test for action $c$ at information set II
1: Construct the sequence-form matrices $(\mathbf{A},\mathbf{E},\mathbf{F})$ for game $G$
2: Let $i_{c}$ denote the sequence index ending in action $c$, and let $i_{p}$ denote the parent sequence
3: Compute $S_{I}$, the set of opponent sequences that reach information set II
4: Solve LP (5) and obtain value $v_{5}$
5: Solve LP (6) and obtain value $v_{6}$
6:if $v_{5}>v_{6}$ then
7:  return action $c$ is strictly dominated
8:else if $v_{5}<v_{6}$ then
9:  return action $c$ is not dominated
10:end if
11: Solve LP (7) and obtain value $v_{7}$
12: Solve LP (8) and obtain value $v_{8}$
13:if $v_{7}>v_{8}$ then
14:  return action $c$ is weakly dominated
15:else
16:  return action $c$ is not dominated
17:end if
@@
There exists a polynomial-time algorithm that determines whether a given action is strictly dominated in a two-player perfect-recall extensive-form game of imperfect information with publicly observable actions.
@@
There exists a polynomial-time algorithm that determines whether a given action is weakly dominated in a two-player perfect-recall extensive-form game of imperfect information with publicly observable actions.
@@
Iterated removal of strictly and weakly dominated actions can be performed in polynomial time in a two-player perfect-recall extensive-form game of imperfect information with publicly observable actions.